OpenStudy (anonymous):

Find a point on the graph of y=4x whose distance to the origin is 2.

OpenStudy (anonymous):

What does it mean by origin?

OpenStudy (anonymous):

The point (0,0)

OpenStudy (anonymous):

so how would I set this up to solve?

OpenStudy (anonymous):

\[\sqrt{(x^2+y^2)}=2;y=4x\]

OpenStudy (anonymous):

\[x^2+(4x)^2=4;\]

OpenStudy (anonymous):

so then I would have \[x ^{2}+16x-4=0 \] and then use quadratic?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

that can't be right b/c I have 2 x^2 then

OpenStudy (anonymous):

so is it \[17x ^{2}-4\]

OpenStudy (anonymous):

\[17x^2=4; x=\pm2/\sqrt17\]

OpenStudy (anonymous):

Do the calculation cautiously.

OpenStudy (anonymous):

My answers choices are (1,1) \[(2\sqrt{17}/17, 8\sqrt{17/17})\], none of the above , and (2,0). I went with none of the above. Is this right?

OpenStudy (anonymous):

I feel it is because none of the other numbers matched.

OpenStudy (anonymous):

Check you problem and see whether it is right.

OpenStudy (anonymous):

I did I was just asking your opinion.

OpenStudy (anonymous):

I already turned it it.

OpenStudy (anonymous):

\[\frac{2}{\sqrt{17}}=\frac{2\sqrt{17}}{17}\]

OpenStudy (anonymous):

and four times that is \[\frac{8\sqrt{17}}{17}\] so that answer would be correct